OpenStudy (anonymous):

Determine the constant c that makes f continuous on (-inifinity, infinity). f(x) = c^2+sin(πx) if x <2 f(x) = cx^2-4 if x ≥2

OpenStudy (anonymous):

anyone? I really need help. I haven't been able to figure it out, my answers don't match the correct answer

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (mayankdevnani):

diificult

OpenStudy (mayankdevnani):

put f(4)

OpenStudy (anonymous):

it's supposed to be c=2 but i don't get how to get there

OpenStudy (anonymous):

are those the pieces of a piece-wise function? if yes, why are both defined for x<2 ?

OpenStudy (anonymous):

oh sorry! i will correct it

OpenStudy (anonymous):

since the two pieces differently to the left and to the right of x=2, the y-values must be the same at x=2: \(\Large c^2+sin(\pi \cdot x)=c \cdot x^2-4 \) \(\Large c^2+sin(\pi \cdot 2)=c \cdot 2^2-4 \) \(\Large c^2=4c-4 \) \(\Large c^2-4c+4=0 \) c = 2

OpenStudy (anonymous):

sorry, i meant to type: since the two pieces are defined differently to the left and right of x=2, ....

OpenStudy (anonymous):

thank you so much!